OpenStudy (anonymous):

Thevolumer V of a cone (V=1/3pir^2h) is increasing at the rate of 28pi cubic units per second. At the instant when the radius r of the cone is 3 units, its volume is 12pi cubic units and the radius is increasing at 1/2 unit per second…. 1) at the instant when the radius of the cone is 3 units, what is the rate of change of the area of the base? PLEASE HELP….

OpenStudy (anonymous):

\[V(t)=\frac{1}{3}r^2(t)h(t)\] \[V'(t)=\frac{1}{3}\left(2r(t)r'(t)h(t)+r^2(t)h'(t)\right)\] is a start

OpenStudy (anonymous):

that being the product rule (plus chain rule) you are told \(V'=28,r=3,r'=\frac{1}{2}\)

OpenStudy (anonymous):

you have to solve for \(h'\) and you are done

OpenStudy (anonymous):

well i guess you also have to solve for \(h\) but you are given enough information for that too

OpenStudy (anonymous):

how did you get the first equation?@satellite73